# Faithful Logic Embeddings in HOL – A recipe to have it all: deep and shallow, automated and interactive, heavy and light, proofs and counterexamples, meta and object level

### v1

@@ 7 Faithfulness automated
Figure 2: Proving faithfulness between different the introduced deep and shallow embeddings of PML in HOL by automated induction.
@@ 8 Experiments
To provide an idea of the different possibilities that are now possible with the linked embeddings, in particular also for logics education, some experiments are reported in this section. The experiments performed also test the soundness of our embeddings, and assess proof automation and counterexample finding for them. The embedding that has performed best in terms of proof automation with sledgehammer and counterexample finding with nitpick is the minimal shallow embedding. This is the reason why file PMLinHOL_shallow_minimal_tests.thy, as shown in Fig. 3 is presented here. The corresponding experiment files with the deep embedding and the maximal shallow embedding are shown in Appendix 0.B.
@@
The first experiments (see lines 4-7) confirm that the schematic axioms H1-H3 and the modus ponens rule MP of the propositional Hilbert calculus from §2 are implied by the minimal shallow embedding. The auto tactic, and with it, the simplifier simp are used, which has access to the definitions in bundle DefM. The $\varphi$, $\psi$ and $\gamma$ are schematic variables of type $\sigma$ in the metalogic HOL. The proofs were originally found by calling sledgehammer.
@@
The next experiment (in lines 9-14) demonstrates automated and interactive reasoning using the schematic axioms and rules of the Hilbert calculus introduced earlier. The proof problem is to show that the formula schema $\varphi\supset^{m}\varphi$ is actually derivable in the given Hilbert calculus. By preventing access to the definitions in the bag DefM, e.g. with the command sledgehammer(H1 H2 H3 MP), the automated theorem provers orchestrated by sledgehammer can be forced to search for proofs in the given Hilbert calculus only, and they succeed here. They find out that only the schematic axioms H1 and H2 and modus ponens MP are needed to obtain the result and return this information to the user. In the given situation reconstruction of this result is then possible with Isabelle’s trusted metis tool. For educational purposes it is relevant that proofs from the literature can be easily mapped one to one to proofs in our setting. This is illustrated here with the provision of an interactive proof (in lines 10-14) which is following a proof from a Wikipedia entry [56] on Hilbert systems. Further derived axiom schemata are proved automatically at the level of the Hilbert calculus (in lines 15 and 16). Of course, such attempts can also fail (as was the case in line 17), and then either interactive proof is needed or, as demonstrated here with the call of auto, using the definitions of the shallow embedding in metalogic level HOL may in fact help to still find a proof automatically. It is relevant to remark, that Isabelle/HOL’s proof tactic support and, in particular, its Eisbach package can be used to leverage proof automation on the level of an embedded logic much further. [^6] For an example on how this can be achieved for a highly non-trivial shallow embedding in HOL see [35, 36]. In this work the Eisbach package of Isabelle/HOL was used to define proof methods for an embedding of abstract object theory AOT, which is the logic underlying Zalta’s Principia-Logico Metaphysica [57]. This work also includes the provision of a resolution prover for AOT.
@@
Student exercises in modal logic textbooks typically include exercises on the well known Sahlqvist correspondences [48] (see lines 22-26). Using the definitions in bag DefM, such correspondences are here automated with the theorem provers provided in Isabelle/HOL. Moreover, more exhaustive experiments on such correspondences, also in more ambitious logic settings, have been reported in prior work [6, 37, 45].
Disproving correspondence and other conjectures.
Even more interesting than automatically proving the above correspondences is disproving falsely assumed correspondences. For example, when calling the model finder nitpick to the false conjecture, that in PML models with reflexive accessibility relations $R$ the axiom schema $\Box^{m}\varphi\supset^{m}\Box^{m}\Box^{m}\varphi$ is implied (see line 28), the following counterexample is reported back by it:
@@
As prior work demonstrates, see e.g. [6, 37, 45] and also [11], nitpick regularly provides intuitive and sometimes even surprising counterexamples when used in conjunction with shallow embeddings. [^7] Prior work [6, 37] also illustrates how the counter-examples reported by nitpick could be systematically converted into statements that verifiably disprove the original conjecture in Isabelle/HOL; the pattern-based recipes presented there could indeed be fully automated. This is not only very useful in computer-supported logic education, but has also provided valuable feedback for research studies as reported, e.g. in [11]. However, some non-classical logics have only infinite models, and then the finite model finder nitpick is of no help; this is e.g. the case for Lorini’s T-STIT logic [40], which was shallowly embedded by Meder [41]. The development of infinite model finding tools and their integration into proof assistants such as Isabelle/HOL is thus an important challenge for further research, as is the automated, domain- and application-adapted provision of intuitive counterexample diagrams to the user.
For further false conjectures (in lines 30-32), including modal collapse $\varphi\supset^{m}\Box^{m}\varphi$, intuitive counterexamples are generated by nitpick. Finally, examples of simple modal theorems (in lines 34-35), such as the KB theorem $\Diamond^{m}\Box^{m}\varphi\supset^{m}\Box^{m}\Diamond^{m}\varphi$, are automated efficiently using smt solvers in combination with definition unfolding in metalogic HOL.
@@
Figure 4 presents further proof experiments. After proving the schematic modal principle $K\Diamond$ (in line 5) an object level proof example [49, Example 6.10] is studied (in lines 8-16): $\Box^{m}P^{m}\supset^{m}(\Diamond^{m}Q^{m}\supset^{m}\Diamond^{m}(P^{m}\wedge^{m}Q^{m}))$, where $P$ and $Q$ are now propositional constant symbols in the PML signature $\mathcal{S}$. This statement is again effectively automated via definition unfolding at the metalevel HOL (line 8), while the presented interactive proof again uses the propositional Hilbert calculus from before in combination with the modal rule of necessitation, the distribution axiom of base modal logic $\mathbf{K}$ and the newly introduced $K\Diamond$-principle.
@@
A first observation, interesting for education and research, is that via definition unfolding the schematic axioms are automatically converted into corresponding formulas in metalogic HOL. For the Löb-axiom (in line 4) we thus obtain after definition unfolding the HOL formula $\forall\varphi\,w.\,(\forall v.\,\texttt{R}wv\longrightarrow(\forall z.\,\texttt{R}vz\longrightarrow\varphi z)\longrightarrow\varphi v)\longrightarrow(\forall v.\,\texttt{R}wv\longrightarrow\varphi v)$, as can be seen in the lower white window area of Isabelle/HOL displayed in Fig. 5. When using the maximal shallow embedding instead, the Löb-axiom unfolds into the expected much heavier HOL formula $\forall\varphi\,WR\,V\,x.\,W\,x\longrightarrow(\forall z.\,W\,z\longrightarrow Rxz\longrightarrow(\forall x.\,Wx\longrightarrow Rzx\longrightarrow\varphi WRVx)\longrightarrow\varphi WRVz)\longrightarrow(\forall z.\,Wz\longrightarrow Rxz\longrightarrow\varphi WRVz)$, which also explains why automated reasoning is generally much harder for maximal shallow embeddings than for minimal shallow embeddings. Using such features of the presented framework, students are given the opportunity to e.g. explore frame properties for non-trivial properties on their own, to eventually simplify them and, using proof automation or interactive proof, to show them equivalent to already known properties, as also attempted in Fig. 5 (in lines 5-9). As shown, such proof automation attempts succeed for the following conjectures: (line 5) the converse well-foundedness of the accessibility relation R in combination with the transitivity of R implies the Löb axiom, (line 7) the Löb axiom implies converse well-foundedness of R, and (line 9) the Löb axiom implies the irreflexivity of R. Whenever transitivity is an implicant to be proved, proof automation still fails (see lines 6 and 8). Here interactive proof is needed or, alternatively, improved HOL theorem proving systems, in particular HOL provers that can suitably guess non-trivial instantiations for higher-order order variables as (apparently) required here. [^8] For more on this challenge topic see e.g. the experiments reported in [12].
Performance.
@@
The performance of sledgehammer and nitpick for the different embeddings is briefly discussed: [^9] Of course, due to the relative simplicity of PML and the majority of test examples presented, these experiments are not yet conclusive. However, they provide evidence for the rationally expected result: The minimal shallow embedding performs better than the maximal shallow embedding and the deep embedding. Minimal shallow embedding: Proofs for 36 out of 38 valid statements (theorems, lemmas and substatements in interactive proofs) were found by calls to sledgehammer. [^10] A MacBook Air (Apple M3, 16GB memory, macOS 15.3.1) and Isabelle 2024 in its standard configuration was used, which sets a 60s prover timeout for sledgehammer. The response time was typically less than a second; except for the Löb tests. Theorems Sound1 and Sound2 were not counted. Nitpick reported counterexamples for 5 out of 5 invalid statements. Reconstruction of sledgehammer proofs with trusted tactics failed in 2 cases. Maximal shallow embedding: Proofs for 34 out of 38 valid statements were found by calls to sledgehammer. Countermodels for 4 out of 5 invalid statements were found by nitpick. Reconstruction of sledgehammer proofs with trusted tactics failed in 4 cases. Deep embedding: Proofs for 33 out of 38 valid statements were found by calls to sledgehammer; in particular, there were no proofs delivered for the Löb examples. Countermodels for 5 out of 5 invalid statements were found by nitpick. Reconstruction of sledgehammer proofs with trusted tactics failed in 2 cases.
@@ 9 Related work and implications for LogiKEy
Much related work on the connections between deep and shallow embeddings can actually be found in the literature, see e.g. [54, 34, 46, 55] and the references therein. However, there seems to be no work reported yet which achieved the automation of faithfulness proofs between deep and (different degrees of) shallow embeddings as demonstrated in this paper.
@@
Exemplary work in this category, which Nipkow called the extensional approach [43], includes a mechanization of HOL reasoning in first-order set theory [31], the techniques presented in [34] to overcome some problems of shallow embedding for reasoning over the syntax of embedded languages, a shallow embedding for a pure type systems in first-order logic [25], and various other works such as [47, 38]; see also the references given in [55]. Moreover, note also that the automation support for various quantified non-classical logics as realized in the HOL theorem prover Leo3 [52, 53] is based on shallow embeddings.
@@
The logic-pluralistic knowledge and reasoning methodology LogiKEy[8], see Fig. 6, leverages higher-order logic (HOL) as a unifying metalogic to enable the modeling of diverse object logics, so far, primarily via using shallow semantic embeddings.
LogiKEy has been used in a number of application domains and contexts where different object logics, such as free logic or modal higher-order logic, for example, have been used as a starting point; see e.g. [11, 13, 45]
This paper illustrates how the LogiKEyapproach may benefit in the future from systematically provided, mutually faithful deep and shallow embeddings of all its supported object logics, so that they be used interchangeably. The technique presented in this paper has also the potential to replace or complement many pen and paper proofs to establish faithfulness results in prior works, such as e.g. given in [9, 3, 10], by verified proofs within the proof assistant. This in turn will be relevant for enabling applications in areas that require highest levels of trust. For this note that a particular application direction of LogiKEy is normative and legal reasoning in the context of AI regulation, see also [39].
Moreover, with respect to the LogiKEy context, this paper also illustrates the conceptual changes that were applied in the transition in previous work from lightweight shallow embeddings of non-classical logics, as used for example for encoding higher-order modal logic [9] or conditional logic [3] (which take into account minimal dependencies on possible worlds only), to more heavyweight shallow embeddings, such as public announcement logic [10], in which additional dependencies (on evaluation domains) had to be maintained. In prior work the explicit modeling of atoms, signatures and valuations was also often avoided.
Most importantly, the presented technique for linking deep and shallow embeddings will be very fruitful for furthering the use of LogiKEy in university-level logic courses, including in particular the author’s courses on universal logical reasoning that have been taught at the Free University of Berlin and now at the University of Bamberg for the last decade.

### v2

@@ 7 Faithfulness automated
Figure 3: Automated faithfulness proofs between different embeddings of PML in HOL.
@@ 8 Experiments
The possibilities that the linked embeddings offer, e.g., for logic education, are illustrated with some experiments. The experiments performed also test the soundness of our embeddings, and assess proof automation and counterexample finding for them with sledgehammer and nitpick. The embedding that has performed best in this regard is the minimal shallow embedding. This is why file PMLinHOL_shallow_minimal_tests.thy, see Fig. 4, is presented here. [^6] Corresponding files for the deep and the maximal shallow embeddings are shown in the appendix [5]. The Isabelle/HOL sources are available at http://logikey.org/tree/master/CoursesAndTutorials/2025-CADE-DeepShallow.
@@
The first experiments (see lines 4-7) confirm that the schematic axioms H1-H3 and the modus ponens rule MP of the propositional Hilbert calculus from §2 are implied by the minimal shallow embedding (local’ versions of H1-H3 and MP rule are also implied; this is not shown here). The auto tactic, and with it, the simplifier simp are used, which has access to the definitions in bundle DefM. The $\varphi$, $\psi$ and $\gamma$ are schematic variables of type $\sigma$ in the metalogic HOL. The proofs were originally found by calling sledgehammer.
@@
The next experiment (in lines 9-14) demonstrates automated and interactive reasoning using the schematic axioms and rules of the Hilbert calculus introduced earlier. It is shown that the formula schema $\varphi\supset^{m}\varphi$ is derivable in the given Hilbert calculus. By preventing access to the definitions in the bag DefM, e.g. with the command sledgehammer(H1 H2 H3 MP), the automated theorem provers orchestrated by sledgehammer can be forced to search for proofs in the given Hilbert calculus only, and they succeed here. They find out that only H1 and H2 and MP are needed to obtain the result. In the given situation reconstruction of a proof from H1 and H2 and MP is then possible with Isabelle’s trusted metis tool. For educational purposes it is relevant that proofs from the literature can be easily mapped one to one to proofs in our setting. This is illustrated here with the provision of an interactive proof (in lines 10-14) which is following a proof from a Wikipedia entry [62] on Hilbert systems. Further derived axiom schemata are proved automatically at the level of the Hilbert calculus (in lines 15-16). Of course, such attempts can also fail (as was the case in line 17), and then either interactive proof is needed or, as demonstrated here with the call of auto, using the definitions of the shallow embedding in metalogic HOL may in fact help to still find a proof automatically. Depending on prior knowledge and teaching objectives, students can be encouraged to provide interactive proofs in exercises rather than using automation tools. It is also important to note that Isabelle/HOL’s proof tactics support, and in particular the Eisbach package, can be used to further enhance the automation of proofs at the level of embedded logic. [^7] The Eisbach package was e.g. used in [40, 41] to define and automate proof methods for a non-trivial shallow embedding of abstract object theory AOT [63] in HOL.
@@
Modal logic textbooks typically include exercises on the well known Sahlqvist correspondences [54] (see lines 22-26). Using the definitions in bag DefM, such correspondences are here automated in Isabelle/HOL. More exhaustive experiments on such (and similar) correspondences have been reported in prior work [3, 8, 42, 50].
@@
Figure 5: Counterexample reported by nitpick in line 28 of Fig. 4
Disproving correspondences.
Even more interesting than automatically proving the above correspondences is disproving falsely assumed correspondences. For example, when nitpick is called on the false conjecture, that in PML models with reflexive accessibility relations $R$ the axiom schema $\Box^{m}\varphi\supset^{m}\Box^{m}\Box^{m}\varphi$ is implied (see line 28), the counterexample shown in Fig. 5 is reported back.
As prior work demonstrates, see e.g. [8, 14, 50], nitpick often provides intuitive and sometimes even surprising counterexamples when used in conjunction with shallow embeddings. [^8] Prior work [8, 42] also illustrates how counterexamples reported by nitpick can be systematically converted into statements that verifiably disprove the false conjectures; the method presented there could indeed be fully automated. This is not only useful for teaching logic, but can also be useful for research studies, see e.g. [14]. However, some non-classical logics have only infinite models, and then the finite model finder nitpick is no longer applicable; this is e.g. the case for Lorini’s T-STIT logic [45], which was embedded in HOL by Meder [46]. Developing infinite model finding tools and integrating them into proof assistants such as Isabelle/HOL is thus an important research challenge. So too is the automated, domain-specific provision of intuitive counterexample diagrams like the one on the left in Fig. 5.
For further false conjectures (in lines 30-32 of Fig. 4), including the modal collapse formula $\varphi\supset^{m}\Box^{m}\varphi$, intuitive counterexamples are generated by nitpick. Finally, examples of simple modal theorems (in lines 34-35), such as the KB theorem $\Diamond^{m}\Box^{m}\varphi\supset^{m}\Box^{m}\Diamond^{m}\varphi$, are automated efficiently using smt solvers in combination with definition unfolding in metalogic HOL.
@@
Figure 6 presents further experiments. After proving the schematic modal principle $K\Diamond$ (in line 5) an object level proof example [55, Ex. 6.10] is studied (in lines 8-16): $\Box^{m}p^{m}\supset^{m}(\Diamond^{m}q^{m}\supset^{m}\Diamond^{m}(p^{m}\wedge^{m}q^{m}))$, where $p$ and $q$ are propositional symbols from the PML signature $\mathcal{S}$. This statement is again effectively automated via definition unfolding at the metalevel HOL (line 8), while the presented interactive proof (in lines 9-16)uses the propositional Hilbert calculus from before in combination with the modal rule of necessitation, the distribution axiom of base logic $\mathbf{K}$ and the newly introduced $K\Diamond$-principle.
@@
A first observation, interesting for education and research, is that via definition unfolding the schematic axioms are automatically converted into corresponding formulas in metalogic HOL. For the Löb-axiom (in line 4) we thus obtain after definition unfolding the HOL formula $\forall\varphi\,w.\,(\forall v.\,\texttt{R}wv\longrightarrow(\forall z.\,\texttt{R}vz\longrightarrow\varphi z)\longrightarrow\varphi v)\longrightarrow(\forall v.\,\texttt{R}wv\longrightarrow\varphi v)$. When using the maximal shallow embedding instead, the Löb-axiom unfolds into the expected much heavier HOL formula $\forall\varphi\,WR\,V\,x.\,W\,x\longrightarrow(\forall z.\,W\,z\longrightarrow Rxz\longrightarrow(\forall x.\,Wx\longrightarrow Rzx\longrightarrow\varphi\,WRVx)\longrightarrow\varphi\,WRVz)\longrightarrow(\forall z.\,Wz\longrightarrow Rxz\longrightarrow\varphi\,WRVz)$, which also explains why automated reasoning is generally much harder for maximal shallow embeddings than for minimal shallow embeddings. The utilisation of the features of the presented framework enables students and researchers to explore frame properties for non-trivial properties independently and to subsequently simplify them. Proof automation or interactive proof can then be employed to investigate their equivalence to already known properties, as illustrated in Fig. 7 (in lines 5-9). As shown, such proof automation attempts succeed for the following conjectures: (line 5) the converse well-foundedness of the accessibility relation R in combination with the transitivity of R implies the Löb axiom, (line 7) the Löb axiom implies converse well-foundedness of R, and (line 9) the Löb axiom implies the irreflexivity of R. Whenever transitivity is an implicant to be proved, proof automation still fails (see lines 6 and 8). Here interactive proof is still required or, alternatively, improved HOL theorem proving systems, in particular provers that can suitably guess non-trivial instantiations for higher-order order variables. [^9] For more on this challenge topic see e.g. the experiments reported in [15].
Performance: examples so far.
@@
The performance of sledgehammer and nitpick for the different embeddings for the example proof problems as discussed in this paper so far is briefly discussed: [^10] A MacBook Pro 16,1 (6-Core Intel Core i7 processor, 2,6 GHz, 6 Cores, 256 KB L2 Cache per Core, 12 MB L3 Cache, 16 GB Memory) and Isabelle 2025 in standard configuration was used, which sets a 30s prover timeout for sledgehammer. Minimal shallow embedding: Proofs for 36 out of 38 valid statements (theorems, lemmas and substatements in interactive proofs) were found by calls to sledgehammer. Nitpick reported counterexamples for 5 out of 5 invalid statements. Reconstruction of sledgehammer proofs with trusted tactics failed in 2 cases. Maximal shallow embedding: Proofs for 34 out of 38 valid statements were foundsledgehammer. Countermodels for 4 out of 5 invalid statements were found by nitpick. Reconstruction of sledgehammer proofs with trusted tactics failed in 4 cases. Deep embedding: Proofs for 32 out of 38 valid statements were found sledgehammer; in particular, there were no proofs delivered for the Löb examples. Countermodels for 5 out of 5 invalid statements were found by nitpick. Reconstruction of proofs failed in 2 cases.
Performance: formula classification.
Further experiments were carried out in order to assess the performance of the different embeddings more rigorously. These tests were about the classification of examples of PML formulas with respect to the logics in the modal logic cube [33] defined by the schematic axioms $T:=\Box\varphi\supset\varphi$, $B:=\varphi\supset\Box\Diamond\varphi$ and $4:=\Box\varphi\supset\Box\Box\varphi$, respectively by their semantic counterparts $Ts:=\texttt{reflexiv}\,R$, $Bs:=\texttt{symmetric}\,R$ and $4s:=\texttt{transitiv}\,R$ (with $R$ being the accessibility relation). The example formulas studied were:
$F_{1}{:}&\Diamond\Diamond\varphi\supset\Diamond\varphi&F_{6}{:}&((\Box(\varphi\supset\psi))\wedge\Diamond\Box\neg\psi)\supset\neg\Diamond\psi\\ F_{2}{:}&\Diamond\Box\varphi\supset\Box\Diamond\varphi&F_{7}{:}&\Diamond\varphi\supset\Box(\varphi\vee\Diamond\varphi)\\ F_{3}{:}&\Diamond\Box\varphi\supset\Box\varphi&F_{8}{:}&\Diamond(\Box\varphi)\supset(\varphi\vee\Diamond\varphi)\\ F_{4}{:}&\Box\Diamond\Box\Diamond\varphi\supset\Box\Diamond\varphi&F_{9}{:}&(\Box\Diamond\varphi\wedge\Box\Diamond\neg\varphi)\supset\Diamond\Diamond\varphi\\ F_{5}{:}&\Diamond(\varphi\wedge\Diamond\psi))\supset(\Diamond\varphi\wedge\Diamond\psi)&F_{10}{:}&(\Box(\varphi\supset\Box\psi)\wedge\Box\Diamond\neg\psi)\supset\neg\Box\psi\end{array}$
The automatically identified weakest logics in the modal cube in which these formulas are valid are as follows: $F_{1}{:}$ K4, $F_{2}{:}$ KB, $F_{3}{:}$ KB4, $F_{4}{:}$ KB/K4, $F_{5}{:}$ K4, $F_{6}{:}$ KB4, $F_{7}{:}$ KB4, $F_{8}{:}$ KT/KB, $F_{9}{:}$ KT, and $F_{10}{:}$ KT.
Explained in abstract terms, for all $M\subseteq\{T,B,4\}$, $N\subseteq\{Ts,Bs,4s\}$ and $F\in\{F_{1},\ldots,F_{10}\}$, sledgehammer and nitpick were called to assess the validity of the following sets of proof problems (each one consisting of $803^{2}\times 10=$ 80 entries): $M\models^{d}F$, $N\models^{d}F$, $M\models^{s}F$, $N\models^{s}F$, $M\models^{m}F$, $N\models^{m}F$. sledgehammer and nitpick were thereby invoked in two modes, with and without prior application of Isabelle/HOL’s simplifier simp to the proof problem under consideration. Consequently, for each set of problems, a total of 160 calls to each tool were conducted (80 with prior simplification and 80 without).
The results of these experiments, conducted with the same hardware as mentioned in Footnote 10, are presented in Fig. 8.
nitpick sledgehammer
$A$ $B$ $C$ $D$ $E$ score
Problems  counterex.  no counterex.  unknown  proof found  no proof
$M\models^{m}F$ 84(42/42) 72(36/36) 4(2/2) 73(38/35) 87(42/45) 229
$N\models^{m}F$ 84(42/42) 76(38/38) 0(0/0) 76(38/38) 84(42/42) 236
$M\models^{s}F$ 32(16/16) 0(0/0) 128(64/64) 32(24/8) 128(56/72) 64
$N\models^{s}F$ 84(42/42) 0(0/0) 76(38/38) 66(38/28) 94(42/52) 150
$M\models^{d}F$ 16(10/6) 74(0/74) 70(70/0) 33(16/17) 127(64/63) 123
$N\models^{d}F$ 32(8/24) 56(0/56) 72(72/0) 70(38/32) 90(42/48) 158
Figure 8: Performance results of automated PML formula classification task conducted with nitpick and sledgehammer (explanation: the first entry $73(\textbf{38}/35)$ in column $D$, e.g., states that 72 proofs were reported in total by sledgehammer for the problem set $M\models^{m}F$, with 38 proofs found with prior application of the simplifier simp and 35 without; analogous for the other entries).
The best results are highlighted. These are the maxima for the informative cases ‘counterexample found’ and ‘no (finite) counterexample found’ for nitpick, and ‘proof found’ for sledgehammer. For the ’unknown’ and ’no proof’ cases, which don’t provide any useful information, the minima are shown as the best results. A overall score is reported, which was computed by adding the results obtained for the informative cases.
The results obtained are briefly discussed. Perfect experimentation results were obtained for the problem set $N\models^{m}F$, that is, the minimal shallow embedding when used in combination with the semantical frame conditions $Ts$, $Bs$ and $4s$. In this setting nitpick either provided a counterexample for each formula classification task or reported that there was no (finite) counterexample. The latter judgments were then supplemented by proofs from sledgehammer that confirmed the validity of these formulas. The second best performance was achieved with the minimal shallow embedding when used in combination with the axiom schemata $T$, $B$ and $4$. The maximal shallow embedding returned some good results regarding counterexample finding with nitpick and theorem proving with sledgehammer for the problem set $N\models^{s}F$ when prior simplification was applied. However, in this setting, nitpick was unable to identify any formulas with no (finite) countermodels. For the deep embedding the performance results obtained with nitpick regarding counterexample finding significantly dropped in comparison to the minimal shallow embedding. The numbers reported for the deep embedding regarding the identification of formulas with no (finite) countermodels are still good though. [^11] However, some seemingly self-contradictory answers by nitpick in this setting with and without prior simplification for some problems require further consideration. Interestingly, the maximal number of 38 proofs was still obtained for by sledgehammer in the deep embedding when prior simplification was applied. Future work should extend the experiments presented to more complicated test formulas $F_{i}$ and also include the further axiom schemata (and their corresponding semantic conditions) from the modal logic cube [33]. Additionally, it seems useful to replay, now using the embeddings provided in this paper, the metalogical, modal logic cube verification experiments reported in prior work [3, 8].
@@ 9 Related work and implications for LogiKEy
The notions of deep and shallow embeddings appear to originate from the work of Boulton et al. [24], and the connections between the two have been discussed in numerous related papers, see e.g. [60, 39, 52, 61] and the references therein. However, to the best of the author’s knowledge, no previous work has achieved an integration of deep and (different degrees of) shallow embeddings in the form presented, enabling in particular mutual faithfulness proofs between them.
@@
Exemplary work in this category, which Nipkow called the extensional approach [48], includes a mechanization of HOL reasoning in first-order set theory [36], the techniques presented in [39] to overcome some problems of shallow embedding for reasoning over the syntax of embedded languages, a shallow embedding for pure type systems in first-order logic [29], and various other works such as [53, 43]; see also the references given in [61]. Note also that the automation support for various quantified non-classical logics as realized in the HOL theorem prover Leo3 [58, 59] is based on minimal shallow embeddings.
@@
The logic-pluralistic knowledge and reasoning methodology LogiKEy[10], see Fig. 9, leverages higher-order logic (HOL) as a unifying metalogic to enable the modeling of diverse object logics, so far, primarily via using shallow semantic embeddings. LogiKEy has been used in a range of application domains and contexts. Very different kinds of object logics have been been embedded LogiKEy over the past decade, including, for example, free logic [14], modal higher-order logic [16], public announcement logic [13], conditional logic [4] and dyadic deontic logic with a preference-based semantics [50].
This paper shows how the LogiKEyapproach could benefit in the future from providing faithful, deep and shallow embeddings of all object logics under consideration. This would allow the embeddings to be used interchangeably in applications. The technique presented has also the potential to replace or complement pen-and-paper faithfulness proofs in prior works, such as the ones reported in [12, 4, 13], by computer-verified proofs in Isabelle/HOL. This in turn will be relevant for enabling applications in areas that require highest levels of trust. For this note that a particular application direction of LogiKEy is normative and legal reasoning in the context of AI regulation, see also [44, 10].
This paper also illustrates the conceptual modifications made in previous work when moving from lightweight shallow embeddings of non-classical logics, such as those used to encode higher-order modal logic [12] or conditional logic [4], which consider only dependencies on possible worlds, to more heavyweight shallow embeddings, as, for example, required to embed public announcement logic [13]; in the latter work additional dependencies on evaluation domains had to be encoded. Note on the other hand, that previous work has often avoided the explicit modeling of atoms, signatures, and evaluations, which has been essential for the work presented in this paper. Moreover, the presented technique should eventually help to overcome challenges in application contexts that require the combined use of deep and shallow embeddings. The Fatio protocol for multi-agent argumentation [51] is such an example. Most importantly, the presented technique for linking deep and shallow embeddings will be very fruitful for furthering the use of LogiKEy in university-level logic courses, including the author’s courses on universal logical reasoning that have been taught for many years at FU Berlin and, more recently, at the University of Bamberg.